OpenStudy (anonymous):

convert the following numbers in various bases to numbers in base 10: Convert 456 in base 7 to an equivalent number in base 10.

OpenStudy (anonymous):

6 in the ones place 5 in the sevens place 4 in the forty nines place

OpenStudy (anonymous):

\[456_{(7)} = 4*7^2+5*7^1+6*7^0=196 + 35 + 6 = 237_{(10)}\]

OpenStudy (anonymous):

oh thank you